Internal critical angle in core at core-cladding interface of a step index fiber is 78.6°. The refractive index of cladding is 1.45. find acceptance angle


1
Expert's answer
2021-04-19T17:13:19-0400

GIVEN:-


critical angle in core at core-cladding interface of a step index fiber is 78.6° = "\theta _C" ......(1)


 refractive index of cladding = 1.45



To Find :-


acceptance angle


solution :-


acceptance angle = "\boxed{\theta_A=sin^{-1}(N_A)}" .......(A)


NA= Numerical apertuture


And we know that


"\boxed{\N_A= (\mu_1^2-\mu_2^2)}"


"\mu_1 =" cladding rifrective index

"\mu_2=" core refractive index



and


"\boxed{sin\theta_c= {\mu_2 \over\mu_1}}"


From (1)


78.6° = "\theta _C"

"\implies"

"sin76^o = {1.45\over\mu_2}"


"0.97029573= {1.45\over\mu_2}"


"\mu_2={1.45\over 0.97029573


}"


"\mu_2 = 1.49438975682" ...........(2)



"\implies"

and


"\boxed{\N_A= (\mu_1^2-(1.45)^2)}"

from (2)


"\boxed{\N_A= ((1.49438975682)^2-(1.45)^2)}"


"\boxed{N_A=0.129}" ...........(3)




Now


acceptance angle =


"\boxed{\theta_A=sin^{-1}(N_A)}"



Answer:- acceptance angle

"\boxed{\theta_A=7.41181^o}"